# Locally Repairable Convertible Codes: Improved Lower Bound and General Construction

### v1

@@ V-A Super-linear length
where $i_{1}\neq i_{2},i_{1},i_{2}\in[t]$, and $\beta_{i}$ is a coset leader of the coset containing $\cup_{j\in\mathcal{P}_{i}}G_{j}$. Therefore, the $t$ initial codes can be the same one code by Remark 4. If the initial code is with super-linear length $n_{I}$, then the length of the final code is $\Theta_{t}(tn_{I})$ which can be super-linear.
@@ V-B Maximally recoverable codes
In general, the local parity checks make it possible for LRCs to recover some predetermined erasure patterns whose size exceeds the minimum Hamming distance. However, it remains an open question whether LRCCs have a similar property. In this subsection, we apply explicit constructions to address this question. More specifically, we consider the case of maximally recoverable (MR) codes, which can recover from all information-theoretically recoverable erasure patterns.
@@
Let $\mathcal{C}$ be an $[n,k,d]_{q}$ code with $(r,\delta)_{a}$-locality, and define $\mathcal{S}=\{S_{i}:i\in[n]\}$, where $S_{i}$ is an $(r,\delta)$-repair set for coordinate $i$. The code $\mathcal{C}$ is said to be a maximally recoverable (MR) code if $\mathcal{S}$ is a partition of $[n]$, and for any $R_{i}\subseteq S_{i}$ such that $|S_{i}\setminus R_{i}|=\delta-1$, the punctured code $\mathcal{C}|_{\cup_{1\leqslant i\leqslant n}R_{i}}$ is an MDS code. Specially, if each $S_{i}\in\mathcal{S}$ is of size $|S_{i}|=r+\delta-1$, the code $\mathcal{C}$ is said to be an $(n,r,h,\delta,q)$-MR code, where $m=\frac{n}{r+\delta-1}$ and $h=mr-k$.
@@
Base Code B([3, Construction A]): Let $\alpha_{1},\alpha_{2},\dots,\alpha_{r+\delta-1}$ be $r+\delta-1$ distinct elements in $\mathbb{F}_{q}^{*}$. Let
@@
$H=\begin{pmatrix}P&O&\cdots&O\\ O&P&\cdots&O\\ \vdots&\vdots&\ddots&\vdots\\ O&O&\cdots&P\\ D(\gamma^{0},\bm{\beta},h,r+\delta-1)&D(\gamma^{1},\bm{\beta},h,r+\delta-1)&\cdots&D(\gamma^{m-1},\bm{\beta},h,r+\delta-1)\end{pmatrix},$ (27)
@@
Therefore, we can derive the following corollary.
Corollary 3:
Let $q\geqslant\max\{r+\delta,m+1\}$, $m\geqslant v\geqslant\lceil\frac{h}{r}\rceil$, and $i_{1}<i_{2}<\cdots<i_{v}$ where $i_{j}\in[m]$ for $j\in[v]$. Then the code $\mathcal{C}$ given by the parity check matrix $H||_{\{i_{1},i_{2},\dots,i_{v}\}}$ is an $(n=v(r+\delta-1),r,h,\delta,q^{h})$-MR code, where $H$ is given by (27).
Corollary 4:
Let $u$ be a positive number and $q\geqslant\max\{r+\delta,m+1\}$. Let the code $\mathcal{C}$ be from Base Code B satisfying $r\mid h$, i.e., the dimension of $\mathcal{C}$ is a positive integer $k=(m-\frac{h}{r})r$. If $u\mid m-\frac{h}{r}$, then there exists an access-optimal LRCC with parameters
@@
Let $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-h/r}{u}]$, and $\mathcal{P}=\{m-h/r+1,\dots,m\}$. Then $\mathcal{C}_{i}$ given by the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[\frac{m-h/r}{u}]$ are the initial codes, where $H$ is given by (27). According to the proof of Theorem 3, the conversion procedure is (18) and the final code is $\mathcal{C}$ with the parity check matrix $H$. Therefore, there exists an access-optimal LRCC. According to Corollary 3, the initial codes are $((u+h/r)(r+\delta-1),r,h,\delta,q^{h})$-MR codes. ∎
@@
In Corollary 4, $H||_{\mathcal{P}_{i}}=D_{i}\cdot H||_{\mathcal{P}_{1}}$ where
$D_{i}=\begin{pmatrix}I_{u(\delta-1)\times u(\delta-1)}&O_{u(\delta-1)\times 1}&O_{u(\delta-1)\times 1}&\cdots&O_{u(\delta-1)\times 1}\\ O_{1\times u(\delta-1)}&1&0&\cdots&0\\ O_{1\times u(\delta-1)}&0&\prod_{j=0}^{1-1}(\gamma^{i-1})^{q^{j}}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ O_{1\times u(\delta-1)}&0&0&\cdots&\prod_{j=0}^{h-1}(\gamma^{i-1})^{q^{j}}\\ \end{pmatrix}.$
Therefore, the $t$ initial codes can be the same one code by Remark 4 when $h=r$.
@@ VI Conclusion
This paper studies the code conversion of locally repairable codes in the merge regime. We improve the lower bound on access cost which is relevant to $(r,\delta)$. As a consequence, we give a general construction of access-optimal LRCCs. Then, based on the general construction, we show an explicit construction of access-optimal LRCCs such that the initial and final codes are with super-linear length. Also, the initial and final codes can be required with maximally recoverable property.

### v2

@@ V-A Super-linear length
where $i_{1}\neq i_{2},i_{1},i_{2}\in[t]$, and $\beta_{i}$ is a coset leader of the coset containing $\cup_{j\in\mathcal{P}_{i}}G_{j}$. Namely, the parity check matrix $H_{i}$ of the initial code $\mathcal{C}_{i}$ for $i\in[t]$ in Theorem 4 is
$H_{i}$ $=\begin{pmatrix}U_{i,1}&\cdots&\mathbf{0}\\ \vdots&\ddots&\vdots\\ \mathbf{0}&\cdots&U_{i,u+1}\\ V_{i,1}&\cdots&V_{i,u+1}\end{pmatrix}$
$=\begin{pmatrix}1&\cdots&0&\cdots&0&\cdots&0&0&\cdots&0\\ \vdots&\ddots&\vdots&&\vdots&&\vdots&\vdots&&\vdots\\ 0&\cdots&\beta_{i}^{\delta-2}&\cdots&0&\cdots&0&0&\cdots&0\\ \vdots&&\vdots&\ddots&\vdots&&\vdots&\vdots&&\vdots\\ 0&\cdots&0&\cdots&1&\cdots&0&0&\cdots&0\\ \vdots&&\vdots&&\vdots&\ddots&\vdots&\vdots&&\vdots\\ 0&\cdots&0&\cdots&0&\cdots&\beta_{i}^{\delta-2}&0&\cdots&0\\ 0&\cdots&0&\cdots&0&\cdots&0&\beta_{i}^{\delta-1}&\cdots&0\\ \vdots&&\vdots&&\vdots&&\vdots&\vdots&\ddots&\vdots\\ 0&\cdots&0&\cdots&0&\cdots&0&0&\cdots&\beta_{i}^{d-2}\end{pmatrix}\begin{pmatrix}1&\cdots&1&\cdots&0&\cdots&0\\ \vdots&\ddots&\vdots&&\vdots&&\vdots\\ \alpha_{1,1}^{\delta-2}&\cdots&\alpha_{1,r+\delta-1}^{\delta-2}&\cdots&0&\cdots&0\\ \vdots&&\vdots&&\vdots&&\vdots\\ 0&\cdots&0&\cdots&1&\cdots&1\\ \vdots&&\vdots&&\vdots&\ddots&\vdots\\ 0&\cdots&0&\cdots&\alpha_{u+1,1}^{\delta-2}&\cdots&\alpha_{u+1,r+\delta-1}^{\delta-2}\\ \alpha_{1,1}^{\delta-1}&\cdots&\alpha_{1,r+\delta-1}^{\delta-1}&\cdots&\alpha_{u+1,1}^{\delta-1}&\cdots&\alpha_{u+1,r+\delta-1}^{\delta-1}\\ \vdots&&\vdots&&\vdots&&\vdots\\ \alpha_{1,1}^{d-2}&\cdots&\alpha_{1,r+\delta-1}^{d-2}&\cdots&\alpha_{u+1,1}^{d-2}&\cdots&\alpha_{u+1,r+\delta-1}^{d-2}\\ \end{pmatrix}$,
where $\cup_{j\in[u+1]}\{\alpha_{j,1},\dots,\alpha_{j,r+\delta-1}\}$ is a multiplicative group of $\mathbb{F}_{q}^{*}$ and $\beta_{i}$ is a coset leader for $i\in[t]$. Therefore, the $t$ initial codes can be the same code by Remark 4. If the initial code has super-linear length $n_{I}$, then the length of the final code generated by Theorem 4 is $\Theta_{t}(tn_{I})$ which can be super-linear.
Example 1:
We give an example for access-optimal $(8,1)_{7^{2}}$ LRCC with parameters $u=2,r=2,\delta=2,d=4,m=17$ in Theorem 4. Since $x^{2}-3$ is an irreducible polynomial over $\mathbb{F}_{7}$, let $\mathbb{F}_{7^{2}}=\mathbb{F}_{7}[x]/(x^{2}-3)=\mathbb{F}_{7}(\beta)$, where $\beta^{2}-3=0$. Then, $\{1,\beta\}$ is a basis of $\mathbb{F}_{7^{2}}$ over $\mathbb{F}_{7}$. Clearly, $\mathbb{F}_{7}^{*}$ is a multiplicative subgroup of $\mathbb{F}_{7^{2}}^{*}$. There are $8$ cosets of $\mathbb{F}_{7}^{*}$, i.e., $\mathbb{F}_{7}^{*}$ and $(i+\beta)\mathbb{F}_{7}^{*},i=0,1,\dots,6$. Let $G_{0}=\{5,6,0\},G_{1}=\{1,2,3\},G_{2}=\{3,4,5\}$. Obviously, $G_{0},G_{1},G_{2}$ satisfy the relationship (37). Therefore, the corresponding initial code $\mathcal{C}_{0}$, based on Base Code A and $G_{0},G_{1},G_{2}$, has the code length $9$ which is large than the size of $\mathbb{F}_{7}$. According to Remark 7, let the initial code $\mathcal{C}_{i}$ is given by Base Code A and $(i-1+\beta)G_{0},(i-1+\beta)G_{1},(i-1+\beta)G_{2}$ for $i\in\{1,2,3,\dots,7\}$. Then, the final code $\mathcal{C}_{F}$, based on Base Code A and $G_{0},G_{1},G_{2},(i-1+\beta)G_{1}\cup(i-1+\beta)G_{2},i=1,2,\dots,7$, has the code length $17\times 3=51>49$. It is easy to check that $\mathcal{C}_{F}$ satisfies the requirement (37). Therefore, the initial and final codes are both with super-linear length. Also, the initial codes $\mathcal{C}_{0},\mathcal{C}_{1},\dots,\mathcal{C}_{7}$ are the same one $\mathcal{C}_{0}$.
@@ V-B Maximally recoverable codes
In general, the local parity checks make it possible for LRCs to recover some predetermined erasure patterns whose size exceeds the minimum Hamming distance, e.g., maximally recoverable codes. However, it remains an open question whether LRCCs retain this property. In this subsection, we apply explicit constructions to address this question. More specifically, we consider the case of maximally recoverable codes, which can recover from all information-theoretically recoverable erasure patterns.
@@
Let $\mathcal{C}$ be an $n,k,d]_{q}$ code with $(r,\delta)_{a}$-locality, and define $\mathcal{S}=\{S_{i}:i\in[n]\}$, where $S_{i}$ is an $(r,\delta)$-repair set for coordinate $i$. The code $\mathcal{C}$ is said to be a maximally recoverable (MR) code if $\mathcal{S}$ is a partition of $[n]$, and for any $R_{i}\subseteq S_{i}$ such that $|S_{i}\setminus R_{i}|=\delta-1$, the punctured code $\mathcal{C}|_{\cup_{1\leqslant i\leqslant n}R_{i}}$ is an MDS code. Specifically, if each $S_{i}\in\mathcal{S}$ is of size $|S_{i}|=r+\delta-1$, the code $\mathcal{C}$ is said to be an $(n,r,h,\delta,q)$-MR code, where $m=\frac{n}{r+\delta-1}$ and h=mrkh=mr-k.
@@
Base Code B([4, Construction A]): Let $\alpha_{1},\alpha_{2},\dots,\alpha_{r+\delta-1}$ be $r+\delta-1$ distinct elements in $\mathbb{F}_{q}^{*}$. Let
@@
$H=\begin{pmatrix}P&\bm{0}&\cdots&\bm{0}\\ \bm{0}&P&\cdots&\bm{0}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{0}&\bm{0}&\cdots&P\\ D(\gamma^{0},\bm{\beta},h,r+\delta-1)&D(\gamma^{1},\bm{\beta},h,r+\delta-1)&\cdots&D(\gamma^{m-1},\bm{\beta},h,r+\delta-1)\end{pmatrix}$, (38)
@@
Therefore, we can derive the following theorem.
Theorem 5:
Let $u$ be a positive integer and $q\geqslant\max\{r+\delta,m+1\}$. Let the code $\mathcal{C}$ be a Base Code B satisfying $r\mid h$, i.e., the dimension of $\mathcal{C}$ is a positive integer $k=(m-\frac{h}{r})r$. If $u\mid m-\frac{h}{r}$, then there is an access-optimal LRCC given by Construction A with parameters
@@
Indeed, $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-h/r}{u}]$, and $\mathcal{P}=\{m-h/r+1,\dots,m\}$. This implies that the initial code $\mathcal{C}_{i}$ has parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ and the final code $\mathcal{C}_{F}$ is by means of the parity check matrix $H$, where $H$ is given by (38). Based on Lemma 12 and $r\mid h$, the minimum Hamming distance of $\mathcal{C}$ is $d=\delta+\frac{h}{r}(r+\delta-1)$ which achieves the lower bound (3). According to Theorem 3 where $g=m$ and $\ell=m-h/r$, the LRCC given by Construction A is access-optimal. Noting that $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ also maintain the structure of Base Code B, therefore, by Lemma 12, the initial codes are $((u+h/r)(r+\delta-1),r,h,\delta,q^{h})$-MR codes and the final code is a $(m(r+\delta-1),r,h,\delta,q^{h})$-MR code. ∎
@@
The $t$ initial codes in Theorem 5 could be the distinct codes. In the following case, the initial codes may be the same one. When $H||_{\mathcal{P}_{i}}=D_{i}\cdot H||_{\mathcal{P}_{1}}$ where
$D_{i}=\begin{pmatrix}I_{u(\delta-1)\times u(\delta-1)}&\bm{0}_{u(\delta-1)\times 1}&\bm{0}_{u(\delta-1)\times 1}&\cdots&\bm{0}_{u(\delta-1)\times 1}\\ \bm{0}_{1\times u(\delta-1)}&1&0&\cdots&0\\ \bm{0}_{1\times u(\delta-1)}&0&\prod_{j=0}^{1-1}(\gamma^{i-1})^{q^{j}}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \bm{0}_{1\times u(\delta-1)}&0&0&\cdots&\prod_{j=0}^{h-1}(\gamma^{i-1})^{q^{j}}\\ \end{pmatrix}$,
the $t$ initial codes can be the same code by Remark 4 when $h=r$.
Example 2:
We give an example for access-optimal $(2,1)_{7^{2}}$ LRCC with parameters $u=2,h=2,r=2,\delta=2,m=5$ in Theorem 5. Let $\mathbb{F}_{7^{2}}$ be the finite field in Example 1. It is easy to check that $\gamma=1+\beta$ is a primitive element. Recall that $\{1,\beta\}$ is a basis of $\mathbb{F}_{7^{2}}$ over $\mathbb{F}_{7}$. Let $\alpha_{1}=1,\alpha_{2}=2,\alpha_{3}=3$ and $\beta_{i}=1+\alpha_{i}^{2}\beta$ for $i=1,2,3$. Based on Remark 8, let the initial code $\mathcal{C}_{1}$ be with the parity check matrix
$H_{1}=\begin{pmatrix}1&1&1&0&0&0&0&0&0\\ 0&0&0&1&1&1&0&0&0\\ 0&0&0&0&0&0&1&1&1\\ 1+\beta&1+2^{2}\beta&1+3^{2}\beta&1+\beta&1+2^{2}\beta&1+3^{2}\beta&1+\beta&1+2^{2}\beta&1+3^{2}\beta\\ (1+\beta)^{7}&(1+2^{2}\beta)^{7}&(1+3^{2}\beta)^{7}&(1+\beta)^{7}\gamma&(1+2^{2}\beta)^{7}\gamma&(1+3^{2}\beta)^{7}\gamma&(1+\beta)^{7}\gamma^{2}&(1+2^{2}\beta)^{7}\gamma^{2}&(1+3^{2}\beta)^{7}\gamma^{2}\end{pmatrix}$
and the initial code $\mathcal{C}_{2}$ be with the parity check matrix
$H_{2}=\begin{pmatrix}1&1&1&0&0&0&0&0&0\\ 0&0&0&1&1&1&0&0&0\\ 0&0&0&0&0&0&1&1&1\\ 1+\beta&1+2^{2}\beta&1+3^{2}\beta&1+\beta&1+2^{2}\beta&1+3^{2}\beta&1+\beta&1+2^{2}\beta&1+3^{2}\beta\\ (1+\beta)^{7}\gamma^{2}&(1+2^{2}\beta)^{7}\gamma^{2}&(1+3^{2}\beta)^{7}\gamma^{2}&(1+\beta)^{7}\gamma^{3}&(1+2^{2}\beta)^{7}\gamma^{3}&(1+3^{2}\beta)^{7}\gamma^{3}&(1+\beta)^{7}\gamma^{4}&(1+2^{2}\beta)^{7}\gamma^{4}&(1+3^{2}\beta)^{7}\gamma^{4}\end{pmatrix}$.
Note that $\mathcal{C}_{1}=\mathcal{C}_{2}$ since
$H_{2}=\begin{pmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&\gamma^{2}\end{pmatrix}H_{1}.$
The final code given by the conversion procedure function (35) has the parity check matrix
$H=\left(\begin{array}[]{cccccc|c}1&1&1&0&0&0&\bm{0}_{1\times 9}\\ 0&0&0&1&1&1&\bm{0}_{1\times 9}\\ \hline\cr\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\bm{0}_{3\times 1}&\\ 1+\beta&1+2^{2}\beta&1+3^{2}\beta&1+\beta&1+2^{2}\beta&1+3^{2}\beta&H_{2}\\ (1+\beta)^{7}&(1+2^{2}\beta)^{7}&(1+3^{2}\beta)^{7}&(1+\beta)^{7}\gamma&(1+2^{2}\beta)^{7}\gamma&(1+3^{2}\beta)^{7}\gamma&\end{array}\right),$
where $H_{1}||_{\{1,2\}}=H||_{\{1,2\}},H_{2}||_{\{1,2\}}=H||_{\{3,4\}}$. Clearly, $\mathcal{C}_{F}$ is an MR code.
@@ VI Conclusion
This paper studies the code conversion of locally repairable codes in the merge regime. We improved the lower bound on the access cost which is relevant to $(r,\delta\geqslant 2)$-locality, broadening the previously known result only related to $r$. As a consequence, we proposed a general construction of access-optimal LRCCs. Then, based on the general construction, we showed two explicit constructions of access-optimal LRCCs such that the final code achieves super-linear length when the initial codes have super-linear length, or the final code has maximally recoverable property when the initial codes have maximally recoverable property, for the first time. Notably, our construction represents the first solution applicable to $\delta>2$ scenarios, surpassing previous constructions limited to $\delta=2$.